OpenStudy (anonymous):

Is the system consistent or inconsistent? 2x + y = 7 and x - 2y = 7 consistent inconsistent i dont get it:(

OpenStudy (anonymous):

a because you cant see the slope

OpenStudy (anonymous):

what do you mean the slope is 2x and x

OpenStudy (anonymous):

opps its a because the slope isent parrelle when you make the line chart thing

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20080108094742AA7lL1X This should help you.

OpenStudy (anonymous):

so consistant lines haev the same slope and if it was consistant it would be 2x and 2xs but its different so its inconsistant thanks people a LOT!!

OpenStudy (anonymous):

and this is false? The system, 6x - 4y = 5 and 9x - 6y = 5, is consistent. True False

OpenStudy (anonymous):

False

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

i got it wrong:( Is the system consistent or inconsistent? 2x + y = 7 and x - 2y = 7 was consistant:(

OpenStudy (anonymous):

Did you not look at the link?

OpenStudy (anonymous):

oops yeah i got it mixed up

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

need hulp